im REALLY bad at word problems so im preparing for the cst with some of the old release question and heres one im having trouble turning this into a graph:
A pharmacist mixed some 10%-saline solution
with some 15%-saline solution to obtain 100 mL of a 12%-saline solution. How much of the 10%-saline solution did the pharmacist use in the mixture?
A 60 mL
B 45 mL
C 40 mL
D 25 mL

Answer by mananth(16946)

percent ---------------- quantity
Acid solution I 10----------------x ml
Acid solution II 15----------------100-x ml
Total 12.00% ---------------- 100
10x+15(100-x)=12*100
10x+1500-15x=1200
10x-15x =1200-1500
-5x=-300
/-5
x=60ml 10.00%
40ml15.00%